# DySK-Attn: A Framework for Efficient, Real-Time Knowledge Updating in Large Language Models via Dynamic Sparse Knowledge Attention

### v1

@@ 1 Introduction
Large Language Models (LLMs) have marked a paradigm shift in artificial intelligence, demonstrating remarkable capabilities in natural language understanding, generation, and reasoning across a multitude of domains [1]. Their success in knowledge-intensive tasks has positioned them as foundational tools for applications ranging from sophisticated question answering systems to complex problem-solving agents capable of using external tools [2, 3]. The underlyi ng principle of these models is their ability to internalize vast amounts of factual knowledge from extensive training corpora, effectively creating an implicit knowledge base [4, 5]. However, this very strength introduces a critical vulnerability: the knowledge encoded within their parameters is inherently static, bound to a "knowledge cut-off" date. This static nature renders them unable to reason about events or information that emerged post-training, leading to factual inaccuracies or "hallucinations" when confronted with novel topics [6]. The challenge of keeping these models current is non-trivial, as complete retraining is computationally prohibitive and unsustainable for real-world, dynamic environments, such as those encountered in continuous sensing applications where new data patterns constantly emerge [7]. This need for adaptability is a shared challenge in many AI fields, including the development of robust federated learning systems.
@@
To surmount this limitation, Retrieval-Augmented Generation (RAG) has emerged as a dominant paradigm [1, 8]. RAG-based methods enhance LLMs by dynamically retrieving information from external corpora at inference time. The field has seen rapid advancements, with sophisticated frameworks like Self-RAG, which learns to critique its own retrieved documents [9], and CRAG, which introduces a corrective mechanism to improve retrieval robustness [10]. Other approaches focus on optimizing the retrieval process itself, such as determining the optimal granularity of information to retrieve [11], developing active retrieval strategies [12, 13], or combining retrieval with chain-of-thought reasoning [14]. While powerful, many RAG systems rely on unstructured text, which can be noisy. This has motivated research into leveraging structured knowledge sources, most notably Knowledge Graphs (KGs), which offer a more organized representation of facts [15]. Integrating KGs allows for more precise reasoning, as demonstrated in tasks like visual dialog where context-aware graph inference is crucial [16], and in unified question-answering systems that operate over heterogeneous data formats [17]. Yet, a significant challenge remains in how to efficiently fuse the symbolic knowledge of KGs with the sub-symbolic representations of LLMs without incurring substantial overhead [18].
@@
Furthermore, as LLMs are tasked with more complex reasoning, often modeled as a "graph of thoughts" [19], the efficiency of their core architecture becomes paramount. The quadratic complexity of the standard Transformer attention mechanism poses a bottleneck for processing long contexts [20]. This has spurred the development of more efficient architectures, including sparse attention models like Longformer [20] and Big Bird , Mixture-of-Experts (MoE) models like Switch Transformers [21, 22] and Mixtral [23], and even entirely new paradigms like state-space models (Mamba) [24] and other efficient long-context mechanisms [25]. These architectural innovations, alongside hardware-aware optimizations like FlashAttention [26, 27], are critical for building scalable systems. The challenge of efficiency is not unique to LLMs; it is a central theme in many areas of deep learning, from model compression through quantization to designing lightweight models for gesture recognition on commodity devices .
@@ 2.1 Knowledge-Augmented Language Models
The inherent limitation of static knowledge in LLMs has led to the development of knowledge-augmented models, with RAG being the most prominent approach [1, 8]. The core idea of RAG is to supplement the model’s internal knowledge with information retrieved from an external source at inference time. Recent advancements have focused on enhancing the robustness and intelligence of the retrieval process. For instance, Self-RAG [28] introduces a self-reflection mechanism, enabling the model to decide when retrieval is necessary and to evaluate the quality of retrieved passages. Similarly, Corrective RAG (CRAG) [29] employs a retrieval evaluator to assess document relevance and triggers corrective actions, a principle of robustness that is also critical in domains like secure physical layer authentication [10] and reliable facial expression recognition from noisy labels [30].
Further research has explored making the retrieval process more dynamic and context-aware. FLARE [12] and Active Retrieval Augmented Generation propose forward-looking strategies where the model actively anticipates future content to decide what information to retrieve next. Other works explore self-adaptive frameworks to determine if retrieval is even necessary for a given query [31]. The Chain-of-Note methodology enhances robustness by having the model first synthesize "notes" from retrieved documents before generating a final answer, filtering out irrelevant information [32]. While these methods have greatly advanced the field, they predominantly operate on unstructured text, which can lack the explicit relational information needed for complex, multi-hop reasoning. Our work diverges by focusing on structured knowledge sources to facilitate more precise and efficient knowledge integration.
@@ 2.2 Efficient Transformer Architectures and Beyond
The computational cost of the standard Transformer’s self-attention mechanism is a major obstacle for long-context applications. To address this, a significant body of work has focused on developing more efficient architectures. Sparse attention models represent one major line of research, using fixed or learned sparsity patterns to reduce computational complexity. Notable examples include Longformer [20], which uses a combination of local windowed attention and global attention[33] , and Big Bird [34], which employs a mix of random, windowed, and global tokens. The pursuit of efficiency is a universal challenge in deep learning, motivating efforts in model compression [22] and the design of lean, effective models for specific applications like WiFi-based sensing [3, 35].
Another highly successful approach is the Mixture-of-Experts (MoE) paradigm, which activates only a subset of model parameters ("experts") for each input token [21, 23]. Beyond Transformer-based innovations, entirely new architectures have emerged. Mamba [24], based on structured state-space models (SSMs), achieves linear-time complexity. Similarly, RecurrentGemma [36] revisits recurrent architectures to gain efficiency, while other works propose novel attention mechanisms like Grouped-Query Attention [37] to reduce memory bandwidth during inference. Our work leverages the principles of sparsity, not at the level of token-to-token interactions, but at the level of knowledge selection, allowing a standard LLM architecture to interact efficiently with a vast external knowledge base.
@@ 2.3 Fusing Language Models with Knowledge Graphs
Knowledge Graphs (KGs) provide a rich, structured representation of entities and their relationships, making them an ideal source for augmenting LLMs. A comprehensive survey by Pan et al. [15] outlines the evolving relationship between KGs and LLMs. Early work focused on injecting knowledge by training models on KG-derived objectives or using adapters, such as K-Adapter [38], to infuse factual knowledge without full fine-tuning. More recent approaches aim for a deeper fusion. GreaseLM [18], for example, combines a language model with a Graph Neural Network (GNN) to allow simultaneous reasoning over both text and graph structures. The challenge of reasoning over structured data is not new and has been explored in areas like benchmarking micro-action recognition and temporal filtering in video grounding [39].
Furthermore, LLMs are now being used not just as consumers of KGs but also as creators. Research has shown that LLMs can be prompted to extract entities and relations from text to construct KGs automatically [6]. Frameworks like Graph-ToolFormer [40], TALM [41], and StructGPT [42] empower LLMs to use external graph analysis tools and query structured databases, treating the KG as an interactive component. Graph-RAG [43] specifically adapts the RAG framework to use the graph’s topology to guide retrieval. Our work builds on this line of research but focuses specifically on the challenge of interacting with a dynamic KG, where knowledge is not static but continuously evolving, a scenario common in real-world systems from human-machine interaction [9] to anti-interference activity recognition .
@@ 2.4 Continual Learning and Knowledge Updating
The ability to update knowledge continually without catastrophic forgetting is a long-standing goal in machine learning, and it is especially critical for LLMs. Recent surveys provide a comprehensive overview of the challenges and methods in continual learning for LLMs [44, 45]. One approach is model editing, which aims to directly modify specific factual knowledge stored in the model’s parameters. Techniques like ROME [46] and MEMIT [47] have shown promise in locating and altering factual associations. Another direction is knowledge unlearning, which focuses on making a model "forget" specific information [48]. This is particularly relevant for privacy-sensitive applications, a concern that also drives research in areas like preventing keystroke snooping [49] and secure handwriting recognition [50].
Other works focus on architectural solutions to handle data streams, such as StreamingLLM [51], or equipping models with explicit long-term memory [52]. The concept of time-aware language models, which explicitly model the temporal validity of facts, is also highly relevant [53]. This is akin to challenges in audio-visual event analysis where temporal context is key [54]. Some works combine these ideas, proposing retrieval-based approaches for lifelong learning [55]. Our approach addresses the knowledge updating problem not by directly editing the model’s weights, but by externalizing the dynamic knowledge into a KG. This sidesteps the risk of catastrophic forgetting and allows for instantaneous updates. The LLM’s role shifts from being a static repository to being a dynamic reasoner that leverages a constantly fresh external knowledge source, a flexible architecture that resembles the goals of multi-modal emotion recognition systems [56] and intelligent agents that can reason and act [57].
@@ 3.2.1 Knowledge Representation
To enable interaction with the neural model, all entities and relations are projected into a low-dimensional continuous vector space. We employ a rotational KG embedding model, such as RotatE, known for its ability to model complex relations [58]. Each entity $e\in\mathcal{E}$ is represented by a vector $\mathbf{e}\in\mathbb{C}^{d_{k}}$, and each relation $r\in\mathcal{R}$ is represented as a rotation in the complex plane, $\mathbf{r}\in\mathbb{C}^{d_{k}}$ with $|\mathbf{r}_{i}|=1$ for each dimension $i$. The plausibility of a triplet $(h,r,t)$ is measured by a scoring function:
@@ 3.2.2 Real-time Update Interface
A key feature of our KG module is its dynamism. We design a high-throughput, low-latency API that allows for real-time updates to the graph. New factual triplets can be ingested and immediately reflected in the graph structure and its vector representations. When a new triplet $(h,r,t)$ is added, its embedding is initialized and the graph index is updated. To maintain embedding quality over time, we employ a periodic retraining schedule that fine-tunes the embeddings on recent data, a process far more efficient than retraining an entire LLM. This ensures the knowledge base remains current, a critical requirement for applications that must adapt to changing environments, such as those involving human-computer interaction [9] or dynamic facial expression analysis [5].
@@ 3.3 Coarse-Grained Knowledge Retrieval
We use an efficient Approximate Nearest Neighbor (ANN) index to retrieve the top-$N$ entities with the highest scores. These entities, along with their immediate neighbors in the KG, form the candidate subgraph $\mathcal{G}_{sub}$, which is passed to the next stage. This two-step process is crucial for scalability, a lesson learned from large-scale sensing systems that must efficiently parse vast amounts of data [3].
@@ 3.4 Sparse Knowledge Attention Mechanism
1 $\mathbf{K}_{attn},\mathbf{V}_{attn}\leftarrow\text{Linear}(\mathbf{F}),\text{Linear}(\mathbf{F})$;
@@ 2 $\mathbf{Q}_{attn}\leftarrow\text{Linear}(\mathbf{h}_{query})$;
$\mathbf{S}\leftarrow\frac{\mathbf{Q}_{attn}\mathbf{K}_{attn}^{T}}{\sqrt{d_{k}}}$ ;
@@
$\mathcal{I}_{topk}\leftarrow\text{Top-k-Indices}(\mathbf{S},k)$ ;
@@
$\mathbf{A}_{sparse}\leftarrow\text{Softmax}(\mathbf{S}_{sparse})$ ;
@@
3 return $\mathbf{k}_{final}$;
@@ 3.5 Knowledge Fusion and Generation
This injection process allows the external knowledge to directly influence the model’s subsequent computations and final token generation, guiding it towards a more factually accurate and up-to-date response. This method of targeted information injection is conceptually similar to how specialized data, like channel state information in WiFi sensing, is used to enhance model performance for specific tasks like gesture recognition [13].
@@ 4.1.2 Time-Sensitive Question Answering
To evaluate the model’s ability to handle new and evolving knowledge, we use the TemporalWiki dataset [50]. This dataset is specifically designed for temporal question answering, where questions are time-sensitive and require reasoning about facts that change over time. We split the dataset based on the timestamp of the facts. Facts dated before our KG snapshot are considered "seen" knowledge, while facts dated after the snapshot are used to simulate real-time updates. This allows us to directly measure how well a model can assimilate and reason with new information that was not present during its initial training. This setup creates a challenging testbed, reflecting real-world scenarios where systems must remain robust against constantly changing data streams, a challenge also seen in domains like anti-interference activity recognition .
@@ 4.2 Baselines
  • Standard RAG [1]: The classic retrieval-augmented generation model. We implement RAG using a dense retriever (Sentence-BERT) over the textual descriptions of the KG entities. This baseline represents the standard approach to knowledge augmentation using unstructured text.
  • GreaseLM [18]: A representative model for deep fusion of KGs and LLMs. It uses a graph neural network to encode structured information and integrates it with the LLM. This baseline allows us to compare our sparse attention approach with methods that perform dense graph-based reasoning.
  • Model Editing (ROME) [46]: To compare our dynamic update mechanism with an alternative approach, we use ROME to directly edit the LLM’s weights to incorporate new facts. For each new fact from TemporalWiki, we perform a model edit. This baseline highlights the trade-offs between externalizing knowledge versus directly modifying model parameters.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have marked a paradigm shift in artificial intelligence, demonstrating remarkable capabilities in natural language understanding, generation, and reasoning across a multitude of domains lewis2020retrieval. Their success in knowledge-intensive tasks has positioned them as foundational tools for applications ranging from sophisticated question answering systems to complex problem-solving agents capable of using external tools qin2023toolllm, wang2024target. The underlyi ng principle of these models is their ability to internalize vast amounts of factual knowledge from extensive training corpora, effectively creating an implicit knowledge base petroni2019language. However, this very strength introduces a critical vulnerability: the knowledge encoded within their parameters is inherently static, bound to a "knowledge cut-off" date. This static nature renders them unable to reason about events or information that emerged post-training, leading to factual inaccuracies or "hallucinations" when confronted with novel topics ji2023llms. The challenge of keeping these models current is non-trivial, as complete retraining is computationally prohibitive and unsustainable for real-world, dynamic environments, such as those encountered in continuous sensing applications where new data patterns constantly emerge huang2023phaseanti. This need for adaptability is a shared challenge in many AI fields, including the development of robust federated learning systems.
@@
To surmount this limitation, Retrieval-Augmented Generation (RAG) has emerged as a dominant paradigm lewis2020retrieval, guu2020realm. RAG-based methods enhance LLMs by dynamically retrieving information from external corpora at inference time. The field has seen rapid advancements, with sophisticated frameworks like Self-RAG, which learns to critique its own retrieved documents zhang2025wiopen, and CRAG, which introduces a corrective mechanism to improve retrieval robustness ahmed2024crag. Other approaches focus on optimizing the retrieval process itself, such as determining the optimal granularity of information to retrieve sachan2024dense, developing active retrieval strategies zhong2023flare, gu2023attention, or combining retrieval with chain-of-thought reasoning ram2023chainofthought. While powerful, many RAG systems rely on unstructured text, which can be noisy. This has motivated research into leveraging structured knowledge sources, most notably Knowledge Graphs (KGs), which offer a more organized representation of facts pan2024knowledge. Integrating KGs allows for more precise reasoning, as demonstrated in tasks like visual dialog where context-aware graph inference is crucial , and in unified question-answering systems that operate over heterogeneous data formats hui2022unikqa. Yet, a significant challenge remains in how to efficiently fuse the symbolic knowledge of KGs with the sub-symbolic representations of LLMs without incurring substantial overhead zhang2022greaselm.
@@
Furthermore, as LLMs are tasked with more complex reasoning, often modeled as a "graph of thoughts" besta2023graph, the efficiency of their core architecture becomes paramount. The quadratic complexity of the standard Transformer attention mechanism poses a bottleneck for processing long contexts beltagy2020longformer. This has spurred the development of more efficient architectures, including sparse attention models like Longformer beltagy2020longformer and Big Bird , Mixture-of-Experts (MoE) models like Switch Transformers fedus2021switch and Mixtral jiang2024mixtral, and even entirely new paradigms like state-space models (Mamba) gu2023mamba and other efficient long-context mechanisms han2024hyperattention. These architectural innovations, alongside hardware-aware optimizations like FlashAttention dao2022flashattention, dao2023flashattention2, are critical for building scalable systems. The challenge of efficiency is not unique to LLMs; it is a central theme in many areas of deep learning, from model compression through quantization to designing lightweight models for gesture recognition on commodity devices .
@@ 2.1 Knowledge-Augmented Language Models
The inherent limitation of static knowledge in LLMs has led to the development of knowledge-augmented models, with RAG being the most prominent approach lewis2020retrieval, guu2020realm. The core idea of RAG is to supplement the model’s internal knowledge with information retrieved from an external source at inference time. Recent advancements have focused on enhancing the robustness and intelligence of the retrieval process. For instance, Self-RAG huang2024phyfinatt introduces a self-reflection mechanism, enabling the model to decide when retrieval is necessary and to evaluate the quality of retrieved passages. Similarly, Corrective RAG (CRAG) asai2023selfrag employs a retrieval evaluator to assess document relevance and triggers corrective actions, a principle of robustness that is also critical in domains like secure physical layer authentication ahmed2024crag and reliable facial expression recognition from noisy labels zhang2025resup.
Further research has explored making the retrieval process more dynamic and context-aware. FLARE zhong2023flare and Active Retrieval Augmented Generation propose forward-looking strategies where the model actively anticipates future content to decide what information to retrieve next. Other works explore self-adaptive frameworks to determine if retrieval is even necessary for a given query li2024whentoretrieve. The Chain-of-Note methodology enhances robustness by having the model first synthesize "notes" from retrieved documents before generating a final answer, filtering out irrelevant information yu2023chainofnote. While these methods have greatly advanced the field, they predominantly operate on unstructured text, which can lack the explicit relational information needed for complex, multi-hop reasoning. Our work diverges by focusing on structured knowledge sources to facilitate more precise and efficient knowledge integration.
@@ 2.2 Efficient Transformer Architectures and Beyond
The computational cost of the standard Transformer’s self-attention mechanism is a major obstacle for long-context applications. To address this, a significant body of work has focused on developing more efficient architectures. Sparse attention models represent one major line of research, using fixed or learned sparsity patterns to reduce computational complexity. Notable examples include Longformer beltagy2020longformer, which uses a combination of local windowed attention and global attention , and Big Bird zaheer2020big, which employs a mix of random, windowed, and global tokens. The pursuit of efficiency is a universal challenge in deep learning, motivating efforts in model compression and the design of lean, effective models for specific applications like WiFi-based sensing wang2024target, huang2020towards.
Another highly successful approach is the Mixture-of-Experts (MoE) paradigm, which activates only a subset of model parameters ("experts") for each input token fedus2021switch, jiang2024mixtral. Beyond Transformer-based innovations, entirely new architectures have emerged. Mamba gu2023mamba, based on structured state-space models (SSMs), achieves linear-time complexity. Similarly, RecurrentGemma googlerecurrentgemma2024 revisits recurrent architectures to gain efficiency, while other works propose novel attention mechanisms like Grouped-Query Attention ainslie2023gqa to reduce memory bandwidth during inference. Our work leverages the principles of sparsity, not at the level of token-to-token interactions, but at the level of knowledge selection, allowing a standard LLM architecture to interact efficiently with a vast external knowledge base.
@@ 2.3 Fusing Language Models with Knowledge Graphs
Knowledge Graphs (KGs) provide a rich, structured representation of entities and their relationships, making them an ideal source for augmenting LLMs. A comprehensive survey by Pan et al. pan2024knowledge outlines the evolving relationship between KGs and LLMs. Early work focused on injecting knowledge by training models on KG-derived objectives or using adapters, such as K-Adapter wang2021k, to infuse factual knowledge without full fine-tuning. More recent approaches aim for a deeper fusion. GreaseLM zhang2022greaselm, for example, combines a language model with a Graph Neural Network (GNN) to allow simultaneous reasoning over both text and graph structures. The challenge of reasoning over structured data is not new and has been explored in areas like benchmarking micro-action recognition and temporal filtering in video grounding .
Furthermore, LLMs are now being used not just as consumers of KGs but also as creators. Research has shown that LLMs can be prompted to extract entities and relations from text to construct KGs automatically ji2023llms. Frameworks like Graph-ToolFormer liu2023graphtoolformer, TALM parisi2022talm, and StructGPT jiang2023structgpt empower LLMs to use external graph analysis tools and query structured databases, treating the KG as an interactive component. Graph-RAG liu2024graphrag specifically adapts the RAG framework to use the graph’s topology to guide retrieval. Our work builds on this line of research but focuses specifically on the challenge of interacting with a dynamic KG, where knowledge is not static but continuously evolving, a scenario common in real-world systems from human-machine interaction zhang2025wiopen to anti-interference activity recognition .
@@ 2.4 Continual Learning and Knowledge Updating
The ability to update knowledge continually without catastrophic forgetting is a long-standing goal in machine learning, and it is especially critical for LLMs. Recent surveys provide a comprehensive overview of the challenges and methods in continual learning for LLMs wang2024comprehensive, luo2024static. One approach is model editing, which aims to directly modify specific factual knowledge stored in the model’s parameters. Techniques like ROME meng2022locating and MEMIT meng2023massediting have shown promise in locating and altering factual associations. Another direction is knowledge unlearning, which focuses on making a model "forget" specific information jiang2023knowledge. This is particularly relevant for privacy-sensitive applications, a concern that also drives research in areas like preventing keystroke snooping huang2024keystrokesniffer and secure handwriting recognition dhingra2021time.
Other works focus on architectural solutions to handle data streams, such as StreamingLLM xiao2023streamingllm, or equipping models with explicit long-term memory zhong2024memorybank. The concept of time-aware language models, which explicitly model the temporal validity of facts, is also highly relevant feng2025rf. This is akin to challenges in audio-visual event analysis where temporal context is key . Some works combine these ideas, proposing retrieval-based approaches for lifelong learning scialom2024retllm. Our approach addresses the knowledge updating problem not by directly editing the model’s weights, but by externalizing the dynamic knowledge into a KG. This sidesteps the risk of catastrophic forgetting and allows for instantaneous updates. The LLM’s role shifts from being a static repository to being a dynamic reasoner that leverages a constantly fresh external knowledge source, a flexible architecture that resembles the goals of multi-modal emotion recognition systems gu2023wife and intelligent agents that can reason and act yao2022react.
@@ 3.2.1 Knowledge Representation
To enable interaction with the neural model, all entities and relations are projected into a low-dimensional continuous vector space. We employ a rotational KG embedding model, such as RotatE, known for its ability to model complex relations li2022temporal. Each entity $e\in\mathcal{E}$ is represented by a vector $\mathbf{e}\in\mathbb{C}^{d_{k}}$, and each relation $r\in\mathcal{R}$ is represented as a rotation in the complex plane, $\mathbf{r}\in\mathbb{C}^{d_{k}}$ with $|\mathbf{r}_{i}|=1$ for each dimension $i$. The plausibility of a triplet $(h,r,t)$ is measured by a scoring function:
@@ 3.2.2 Real-time Update Interface
A key feature of our KG module is its dynamism. We design a high-throughput, low-latency API that allows for real-time updates to the graph. New factual triplets can be ingested and immediately reflected in the graph structure and its vector representations. When a new triplet $(h,r,t)$ is added, its embedding is initialized and the graph index is updated. To maintain embedding quality over time, we employ a periodic retraining schedule that fine-tunes the embeddings on recent data, a process far more efficient than retraining an entire LLM. This ensures the knowledge base remains current, a critical requirement for applications that must adapt to changing environments, such as those involving human-computer interaction zhang2025wiopen or dynamic facial expression analysis .
@@ 3.3 Coarse-Grained Knowledge Retrieval
We use an efficient Approximate Nearest Neighbor (ANN) index to retrieve the top-$N$ entities with the highest scores. These entities, along with their immediate neighbors in the KG, form the candidate subgraph $\mathcal{G}_{sub}$, which is passed to the next stage. This two-step process is crucial for scalability, a lesson learned from large-scale sensing systems that must efficiently parse vast amounts of data wang2024target.
@@ 3.4 Sparse Knowledge Attention Mechanism
1
[…]
@@ 3 $\mathbf{Q}_{attn}\leftarrow\text{Linear}(\mathbf{h}_{query})$;
4
[…]
@@
5
[…]
@@
6
[…]
@@
7
81exreturn $\mathbf{k}_{final}$;
@@ 3.5 Knowledge Fusion and Generation
This injection process allows the external knowledge to directly influence the model’s subsequent computations and final token generation, guiding it towards a more factually accurate and up-to-date response. This method of targeted information injection is conceptually similar to how specialized data, like channel state information in WiFi sensing, is used to enhance model performance for specific tasks like gesture recognition gu2023attention.
@@ 4.1.2 Time-Sensitive Question Answering
To evaluate the model’s ability to handle new and evolving knowledge, we use the TemporalWiki dataset dhingra2021time. This dataset is specifically designed for temporal question answering, where questions are time-sensitive and require reasoning about facts that change over time. We split the dataset based on the timestamp of the facts. Facts dated before our KG snapshot are considered "seen" knowledge, while facts dated after the snapshot are used to simulate real-time updates. This allows us to directly measure how well a model can assimilate and reason with new information that was not present during its initial training. This setup creates a challenging testbed, reflecting real-world scenarios where systems must remain robust against constantly changing data streams, a challenge also seen in domains like anti-interference activity recognition .
@@ 4.2 Baselines
  • Standard RAG lewis2020retrieval: The classic retrieval-augmented generation model. We implement RAG using a dense retriever (Sentence-BERT) over the textual descriptions of the KG entities. This baseline represents the standard approach to knowledge augmentation using unstructured text.
  • GreaseLM zhang2022greaselm: A representative model for deep fusion of KGs and LLMs. It uses a graph neural network to encode structured information and integrates it with the LLM. This baseline allows us to compare our sparse attention approach with methods that perform dense graph-based reasoning.
  • Model Editing (ROME) meng2022locating: To compare our dynamic update mechanism with an alternative approach, we use ROME to directly edit the LLM’s weights to incorporate new facts. For each new fact from TemporalWiki, we perform a model edit. This baseline highlights the trade-offs between externalizing knowledge versus directly modifying model parameters.